Use 9.2 to evaluate sin π
5
,sin 2π
5
and cos π
5
. (9)
10.2 Let z = cosθ + sin θ.
Then z
n = cos nθ + isin θ for all n ∈ N (by de Moivre) and z
−n = cos nθ − isin nθ.
(a) Show that 2 cos nθ = z
n + z
−n
and 2isin nθ = z
n − z
n
. (2)
(b) Show that 2
n
cosn
θ =
1
Expert's answer
2020-04-24T18:27:56-0400

Let

"\theta=\pi/5, =>5\theta=\pi.\\
3\theta=\pi-2\theta. \\"

Therefore,


"\text{sin}(3\theta)=\text{sin}(\pi-2\theta)=\text{sin}(2\theta)"

if "\theta<90."

Next,


"\text{sin}3\theta=\text{sin}2\theta,\\
3\text{sin}\theta-4\text{sin}^3\theta=2\text{sin}\theta\text{cos}\theta\space||\cdot\frac{1}{\text{sin}\theta},\\
3-4\text{sin}^2\theta=2\text{cos}\theta,\\
3-4(1-\text{cos}^2\theta)=2\text{cos}\theta,\\
4\text{cos}^2\theta-2\text{cos}\theta-1=0."

We notice that this is a quadratic equation where cos theta is like a variable:


"\text{cos}\theta=\frac{1\pm\sqrt5}{4},"

we take only positive root because theta cannot give negative cosine,


"\text{cos}\theta=\text{cos}\frac{\pi}{5}=\frac{1+\sqrt5}{4},\\
\text{sin}\theta=\text{sin}\frac{\pi}{5}=\sqrt{1-\text{cos}^2\theta}=\frac{\sqrt{10-2\sqrt5}}{4}.\\
\text{sin}\frac{2\pi}{5}=\text{sin}(2\theta)=2\text{sin}\theta\text{cos}\theta=\\
=2\cdot\frac{\sqrt{10-2\sqrt5}}{4}\cdot\frac{1+\sqrt5}{4}
=\frac{\sqrt{{10+2\sqrt5}}}{4}."

(a)


"z^n+z^{-n}=(\text{cos}\theta+i\text{sin}\theta)^n+(\text{cos}\theta+i\text{sin}\theta)^{-n},\\"

according to de Moivre theorem this is


"\text{cos}(n\theta)+i\text{sin}(n\theta)+\text{cos}(-n\theta)+i\text{sin}(-n\theta)=\\
=\text{cos}(n\theta)+i\text{sin}(n\theta)+\text{cos}(-n\theta)-i\text{sin}(n\theta)=\\
=2\text{cos}(n\theta)."